OpenStudy (anonymous):

Medal and fan awarded, please help!!! Select the coordinates of two points on the line y = 8. (8, 8) and (8, 0) (8, -8) and (8, 8) (0, 8) and (8, 0) (0, 8) and (8, 8)

Directrix (directrix):

Hint: The two points will have y-coordinates of 8.

Directrix (directrix):

That is because they are on the line y = 8 and all points on that line have y-coordinate 8.

Directrix (directrix):

@CoinBank77 Which two points do you think are correct?

OpenStudy (anonymous):

So, if both points will have 8 as Y, it should be the last one? (0,8) (8,8)

Directrix (directrix):

(0,8) (8,8) Yes, these two points have 8 as the y coordinate. The x coordinates do not matter. The only requirement is that the y be 8.

OpenStudy (anonymous):

Great, thanks again!

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

Mind helping me with two more?

Directrix (directrix):

I will try.

Directrix (directrix):

@CoinBank77 Post

OpenStudy (anonymous):

Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1

OpenStudy (anonymous):